### v3

@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
Furthermore, $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x}):\Omega_{\boldsymbol{m}}\to\mathbb{R}$ is a constant function with on each $\Omega_{\boldsymbol{m},\boldsymbol{i}}$ for $\boldsymbol{i}\in\{0\ldots,K\}^{d}.$ And
(16) $|g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})|\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}})}(4K)^{\frac{d}{p}}$
for all $\boldsymbol{x}\in\Omega_{\boldsymbol{m}}$, where $C_{1}$ and $C_{2}$ are constants independent of $K$.
This proof is similar to that of [14, Lemma C.4.], but we provide detailed proof as follows for readability. Before the proof, we must introduce the partition of unity, average Taylor polynomials, and a lemma.
Definition 4.14 (The partition of unity).
Let $d,K\in\mathbb{N}_{+}$, then
$\Psi=\left\{h_{\boldsymbol{i}}:{\boldsymbol{i}}\in\{0,1,\ldots,K\}^{d}\right\}$
with $h_{\boldsymbol{i}}:\mathbb{R}^{d}\to\mathbb{R}$ for all ${\boldsymbol{i}}\in\{0,1,\ldots,K\}^{d}$ is called the partition of unity $[a,b]^{d}$ if it satisfies
(i): $0\leq h_{\boldsymbol{i}}(\boldsymbol{x})\leq 1$ for every $h_{\boldsymbol{i}}\in\Psi$.
(ii): $\sum_{h_{\boldsymbol{i}}\in\Psi}h_{\boldsymbol{i}}=1$ for every $x\in[a,b]^{d}$.
Definition 4.15.
Let $n\geq 1$ and $f\in W^{n,p}([a,b]^{d})$, $\boldsymbol{x}_{0}\in[a,b]^{d}$ and $r>0$ such that for the ball $B(\boldsymbol{x}_{0}):=B(\boldsymbol{x}_{0})_{r,|\cdot|}$ which is a compact subset of $[a,b]^{d}$. The corresponding Taylor polynomial of order $n$ of $f$ averaged over $B$ is defined for
(17) $Q^{n}f(x):=\int_{B}T^{n}_{\boldsymbol{y}}f(\boldsymbol{x})b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}$
where
$T^{n}_{\boldsymbol{y}}{f}(\boldsymbol{x})$ $:=\sum_{|\boldsymbol{\alpha}|\leq n-1}\frac{1}{\boldsymbol{\alpha}!}D^{\boldsymbol{\alpha}}{f}(\boldsymbol{y})(\boldsymbol{x}-\boldsymbol{y})^{\boldsymbol{\alpha}},$
$b_{r}(\boldsymbol{x})$ :=$:=\begin{cases}\frac{1}{c_{r}}e^{-\left(1-\left(|\boldsymbol{x}-\boldsymbol{x}_{0}|/r\right)^{2}\right)^{-1}},~{}&|\boldsymbol{x}-\boldsymbol{x}_{0}|<r,\\ 0,~{}&|\boldsymbol{x}-\boldsymbol{x}_{0}|\leq r,\end{cases}$
(18) $c_{r}$ $=\int_{\mathbb{R}^{d}}e^{-\left(1-\left(|\boldsymbol{x}-\boldsymbol{x}_{0}|/r\right)^{2}\right)^{-1}}\,\mathrm{d}x.$
Lemma 4.16 ([14]).
Let $n\geq 1$ and $f\in W^{n,p}([a,b]^{d})$, $\boldsymbol{x}_{0}\in\Omega$ and $r>0$ such that for the ball $B:=B_{r,|\cdot|}(\boldsymbol{x}_{0})$ which is a compact subset of $[a,b]^{d}$. The corresponding Taylor polynomial of order $n$ of $f$ averaged over $B$ can be read as
$Q^{n}f(\boldsymbol{x})=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}.$
Furthermore,
(19) $\left|c_{f,\boldsymbol{\alpha}}\right|\leq C_{2}(n,d)r^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}$
where $C_{2}(n,d)$ is a constant independent of $f$.
Proof 4.17.
Based on [14, Lemma B.9.], $Q^{n}f(x)$ can be read as
(20) $Q^{n}f(\boldsymbol{x})=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}$
where
(21) $c_{f,\boldsymbol{\alpha}}=\sum_{|\boldsymbol{\alpha}+\boldsymbol{\beta}|\leq n-1}\frac{1}{(\boldsymbol{\beta}+\boldsymbol{\alpha})!}a_{\boldsymbol{\beta}+\boldsymbol{\alpha}}\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})\boldsymbol{y}^{\boldsymbol{\beta}}b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}$
for $a_{\boldsymbol{\beta}+\boldsymbol{\alpha}}\leq\frac{(\boldsymbol{\alpha}+\boldsymbol{\beta})!}{\boldsymbol{\alpha}!\boldsymbol{\beta}!}$. Note that
$\left|\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})\boldsymbol{y}^{\boldsymbol{\beta}}b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}\right|\leq\left|\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}\right|$
(22) $\leq$ $\|{f}\|_{W^{n-1,p}(B)}\|b_{r}(x)\|_{L^{q}(B)}\leq cr^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}.$
Then
(23) $\left|c_{f,\boldsymbol{\alpha}}\right|\leq C_{2}(n,d)r^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}$
where $C_{2}(n,d)=\sum_{|\boldsymbol{\alpha}+\boldsymbol{\beta}|\leq n-1}\frac{c}{\boldsymbol{\alpha}!\boldsymbol{\beta}!}$.
The proof of Proposition 4.13 is based on average Taylor polynomials and the Bramble–Hilbert Lemma [4, Lemma 4.3.8].
Definition 4.18.
Let $\Omega,~{}B\in\mathbb{R}^{d}$. Then $\Omega$ is called stared-shaped with respect to $B$ if
$\overline{\text{conv}}\left(\{\boldsymbol{x}\}\cup B\subset\Omega\right),~{}\text{for all }\boldsymbol{x}\in\Omega.$
Definition 4.19.
Let $\Omega\in\mathbb{R}^{d}$ be bounded, and define
:=$\mathcal{R}:=\left\{r>0:\begin{array}[]{l}\text{ there exists }\boldsymbol{x}_{0}\in\Omega\text{ such that }\Omega\text{ is }\\ \text{ star-shaped with respect to }B_{r,|\cdot|}\left(\boldsymbol{x}_{0}\right)\end{array}\right\}.$
Then we define
$r_{\max}^{\star}:=\sup\mathcal{R}\quad\text{ and call }\quad\gamma:=\frac{\operatorname{diam}(\Omega)}{r_{\max}^{\star}}$
the chunkiness parameter of $\Omega$ if $\mathcal{R}\not=\emptyset$.
Lemma 4.20 (Bramble–Hilbert Lemma [4, Lemma 4.3.8]).
Let $\Omega\in\mathbb{R}^{d}$ be open and bounded, $\boldsymbol{x}_{0}\in\Omega$ and $r>0$ such that $\Omega$ is the stared-shaped with respect to $B:=B_{r,|\cdot|}\left(\boldsymbol{x}_{0}\right)$, and $r\geq\frac{1}{2}r_{\max}^{\star}$. Moreover, let $n\in\mathbb{N}_{+}$, $1\leq p\leq p$ and denote by $\gamma$ by the chunkiness parameter of $\Omega$. Then there is a constant $C(n,d,\gamma)>0$ such that for all $f\in W^{n,p}(\Omega)$
$\left|f-Q^{n}f\right|_{W^{k,p}(\Omega)}\leq C(n,d,\gamma)h^{n-k}|f|_{W^{n,p}(\Omega)}\quad\text{ for }k=0,1,\ldots,n$
where $Q^{n}f$ denotes the Taylor polynomial of order $n$ of $f$ averaged over $B$ and $h=\operatorname{diam}(\Omega)$.
Proof 4.21 (Proof of Proposition 4.13).
Without loss of generalization, we prove the case for $\boldsymbol{m}=(1,1,\ldots,1)=:\boldsymbol{m}_{*}$.
Denote $E:W^{n,p}([a,b]^{d})\to W^{n,p}(\mathbb{R}^{d})$ be an extension operator [36] and set $\tilde{f}:=Ef$ and $C_{E}$ is the norm of the extension operator.
Define $p_{f,\boldsymbol{i}}$ as the average Taylor polynomial Definition 4.15 in
$B_{\boldsymbol{i},K}:=B_{\frac{1}{4K}\cdot(b-a),|\cdot|}\left(a+\frac{8\boldsymbol{i}+3}{8K}\cdot(b-a)\right)$
i.e.
(24) $p_{f,\boldsymbol{i}}:=\int_{B_{\boldsymbol{i},K}}T^{n}_{\boldsymbol{y}}\tilde{f}(\boldsymbol{x})b_{\frac{1}{4K}}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}.$
Based on Lemma 4.16, $p_{f,\boldsymbol{i}}$ can be read as
(25) $p_{f,\boldsymbol{i}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{i},\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}},$
where
(26) $|c_{f,\boldsymbol{i},\boldsymbol{\alpha}}|\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(B_{\boldsymbol{i},K})}(4K)^{\frac{d}{p}}\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}.$
The reason to define average Taylor polynomial on $B_{\boldsymbol{i},K}$ is to use the Bramble–Hilbert Lemma 4.20 on
$\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}=B_{\frac{3}{8K}\cdot(b-a),\|\cdot\|_{\ell_{\infty}}}\left(a+\frac{8\boldsymbol{i}+3}{8K}\cdot(b-a)\right)=\prod_{j=1}^{d}\left[a+\frac{i_{j}}{K}\cdot(b-a),a+\frac{3+4i_{j}}{4K}\cdot(b-a)\right].$

### v4

@@ 4.3.1 Proof of Theorem 2
Furthermore, $g_{f,\bm{\alpha},\bm{m}}(\bm{x}):\Omega_{\bm{m}}\to\mathbb{R}$ is a constant function on each $\Omega_{\bm{m},\bm{i}}$ for $\bm{i}\in\{0\ldots,K\}^{d},$ and
$|g_{f,\bm{\alpha},\bm{m}}(\bm{x})|\leq C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}})}(4K)^{\frac{d}{p}}$ (4.14)
for all $\bm{x}\in\Omega_{\bm{m}}$, where $C_{1}$ and $C_{2}$ are constants independent of $K$.
This proof is similar to that of [21, Lemma C.4] and [59, Theorem 8], so we omit detailed proof in this paper.
Remark 3.
The constant $C_{1}$ in Proposition 7 depends linearly on $\max\{|b-a|^{n-i}\}_{i=0}^{2}$. This dependence arises from the Bramble–Hilbert Lemma [6, Lemma 4.3.8]. As shown in the proof of [59, Theorem 8] and [6, Lemma 4.3.8], the approximation constant in the Bramble–Hilbert Lemma scales with $(\operatorname{diam}(\Omega))^{s}$, where $\Omega$ is the approximation domain and $s$ is the difference between the regularity of the function and the order of the norm in which the error is measured. The constant $C_{2}$ depends on the maximal distance between the approximation domain and the origin. Since the target function is approximated locally by polynomials, similar to a Taylor expansion, the coefficients of these polynomials naturally depend on the location within the domain. In our setting, $g_{f,\bm{\alpha},\bm{m}}$ is a piecewise constant function that represents the local coefficients of the polynomial approximation. The bound on $g_{f,\bm{\alpha},\bm{m}}$, and hence the constant $C_{2}$, depends on $\max\{|a|,|b|\}$, and the degree of this dependence is determined by the order of the polynomial. A more detailed analysis can be found in [21, Lemma B.9].
Proof of Theorem 2.
Without loss of the generalization, we consider the case for $\bm{m}_{*}=(1,1,\ldots,1)$. Due to Proposition 7 and setting $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, we have
$\|f-f_{K,\bm{m}_{*}}\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}\leq C_{1}(n,d,|b-a|)\|f(\bm{x})\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2(n-2)/d}L^{-2(n-2)/d}$
$\|f-f_{K,\bm{m}_{*}}\|_{W^{1,p}(\Omega_{\bm{m}_{*}})}\leq C_{1}(n,d,|b-a|)\|f(\bm{x})\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2(n-1)/d}L^{-2(n-1)/d}$
$\|f-f_{K,\bm{m}_{*}}\|_{L^{p}(\Omega_{\bm{m}_{*}})}\leq C_{1}(n,d,|b-a|)\|f(\bm{x})\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2n/d}L^{-2n/d},$ (4.15)
where $f_{K,\bm{m}_{*}}=\sum_{|\bm{\alpha}|\leq n-1}g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})\bm{x}^{\bm{\alpha}}$ for $x\in\Omega_{\bm{m}_{*}}$. Note that $g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})$ is a constant function for $\bm{x}\in\prod_{j=1}^{d}\left[a+\frac{i_{j}}{K}\cdot(b-a),a+\frac{3+4i_{j}}{4K}\cdot(b-a)\right]$ and $\bm{i}=(i_{1},\ldots,i_{d})\in\{0,1,\ldots,K-1\}^{d}$. The remaining part is to approximate $f_{K,\bm{m}_{*}}$ by neural networks.
The way to approximate $g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})$ is similar with [28, Theorem 3.1]. First of all, due to Proposition 5, there is a neural network $\phi_{1}(x)$ with the width $4N+5$ and depth $4L+4$ such that
$\phi(x)=k,x\in\left[a+\frac{k\cdot(b-a)}{K},a+\frac{(k+1)\cdot(b-a)}{K}-\frac{b-a}{4K}\right],~k=0,1,\ldots,K-1.$ (4.16)
Note that we choose $\delta=\frac{1}{4K}\leq\frac{1}{3K}$ in Proposition 5. Then define
$\bm{\phi}_{2}(\bm{x})=\left[\frac{\phi_{1}(x_{1})}{K},\frac{\phi_{1}(x_{2})}{K},\ldots,\frac{\phi_{1}(x_{d})}{K}\right]^{\intercal}.$
For each $p=0,1,\ldots,K^{d}-1$, there is a bijection
$\bm{\eta}(p)=[\eta_{1},\eta_{2},\ldots,\eta_{d}]\in\{0,1,\ldots,K-1\}^{d}$
such that $\sum_{j=1}^{d}\eta_{j}K^{j-1}=p$. Then define
$\xi_{\bm{\alpha},p}=\frac{g_{f,\bm{\alpha},\bm{m}_{*}}\left(\frac{\bm{\eta}(p)}{K}\right)+C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(4K)^{\frac{d}{p}}}{2C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(4K)^{\frac{d}{p}}}\in[0,1],$
where $C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(4K)^{\frac{d}{p}}$ is the bounded of $g_{f,\bm{\alpha},\bm{m}_{*}}$ and $C_{2}$ defined in Proposition 7. Therefore, based on Proposition 6, there is a neural network $\tilde{\phi}_{\bm{\alpha}}(x)$ with the width
$16\left(n+\frac{d}{p}\right)(N+1)\log_{2}(8N)$
and depth $(5L+2)\log_{2}(4L)$ such that $|\tilde{\phi}_{\bm{\alpha}}(p)-\xi_{\bm{\alpha},p}|\leq N^{-2\left(n-\frac{d}{p}\right)}L^{-2\left(n-\frac{d}{p}\right)}$ for $p=0,1,\ldots K^{d}-1$. Denote
$\phi_{\bm{\alpha}}(\bm{x})=C_{2}(n,d,|a|,|b|)(4K)^{\frac{d}{p}}\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}\left[2\tilde{\phi}_{\bm{\alpha}}\left(\sum_{j=1}^{d}\eta_{j}K^{j}\right)-1\right]$
and obtain that
$\left|\phi_{\bm{\alpha}}\left(\frac{\bm{\eta}(p)}{K}\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\frac{\bm{\eta}(p)}{K}\right)\right|$
$=$ $2C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(4K)^{\frac{d}{p}}|\tilde{\phi}_{\bm{\alpha}}(p)-\xi_{\bm{\alpha},p}|$
$\leq$ $4^{\frac{2d+p}{2p}}\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}C_{2}(n,d,|a|,|b|)(NL)^{-2n}.$
Then we obtain that
$\|\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}(\bm{x})\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)\|_{W^{1,p}(\Omega_{\bm{m}_{*}})}=$ $\|\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}(\bm{x})\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)\|_{L^{p}(\Omega_{\bm{m}_{*}})}$
$\leq$ $4^{\frac{2d+p}{2p}}C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(NL)^{-2n}$ (4.17)
which is due to $\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}(\bm{x})\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)$ is a step function, and the first order weak derivative is 0 in $\Omega_{\bm{m}_{*}}$, and we have that
$\|\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}=$ $\|\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)\|_{W^{1,p}(\Omega_{\bm{m}_{*}})}$
$=$ $\|\phi_{\bm{\alpha}}\left(\bm{\phi}_{2}\right)-g_{f,\bm{\alpha},\bm{m}_{*}}\left(\bm{x}\right)\|_{L^{p}(\Omega_{\bm{m}_{*}})}$
$\leq$ $4^{\frac{2d+p}{2p}}C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(NL)^{-2n}.$ (4.18)
Due to (v) in Lemma 3, there is a $\sigma_{2}$ neural network $\phi_{5,\bm{\alpha}}(\bm{x})$ with the width $N$ due to $N>n^{d}$ and depth $\left\lceil\log_{2}L\right\rceil$ due to $L>n$ such that
$\phi_{5,\bm{\alpha}}(\bm{x})=\bm{x}^{\bm{\alpha}},~\bm{x}\in\mathbb{R}^{d}.$ (4.19)
Due to (iv) in Lemma 3, there is a $\sigma_{2}$ neural network $\phi_{6}(\bm{x})$ with the width $4$ and depth $1$ such that
$\phi_{6}(x,y)=xy,~x,y\in\mathbb{R}.$ (4.20)
Now we define the neural network $\gamma_{\bm{m}_{*}}(\bm{x})$ to approximate $f_{K,{\bm{m}_{*}}}(\bm{x})$ in $\Omega_{\bm{m}_{*}}$:
$\gamma_{\bm{m}_{*}}(\bm{x})=\sum_{|\bm{\alpha}|\leq n-1}\phi_{6}\left[\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x})),\phi_{5,\bm{\alpha}}(\bm{x})\right].$ (4.21)
$\gamma_{\bm{m}_{*}}$ is a $\sigma_{2}$-NN with depth $\log_{2}9(L+1)$ and the width $28n^{d+1}(N+d)\log_{2}(8N)$.
The remaining question is to find the error $\mathcal{E}$:
$\widetilde{\mathcal{E}}$:= $\left\|\sum_{|\bm{\alpha}|\leq n-1}\phi_{6}\left[\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x})),\phi_{5,\bm{\alpha}}(\bm{x})\right]-f_{K,\bm{m}_{*}}(\bm{x})\right\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}$
$\leq$ $\sum_{|\bm{\alpha}|\leq n-1}\left\|\phi_{6}\left[\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x})),\phi_{5,\bm{\alpha}}(\bm{x})\right]-g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})\bm{x}^{\bm{\alpha}}\right\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}$
$=$ $\sum_{|\bm{\alpha}|\leq n-1}\left\|\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x}))\bm{x}^{\bm{\alpha}}-g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})\bm{x}^{\bm{\alpha}}\right\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}$
$\leq$ $n^{2}\sum_{|\bm{\alpha}|\leq n-1}\left\|\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x}))-g_{f,\bm{\alpha},\bm{m}_{*}}(\bm{x})\right\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}$
$\leq$ $n^{d+2}4^{\frac{2d+p}{2p}}C_{2}(n,d,|a|,|b|)\|f\|_{W^{n-1,p}(\Omega_{\bm{m}_{*}})}(NL)^{-2n}.$ (4.22)
Combining Eqs. (4.15) and (4.22), we have that there is a DSRN $\gamma_{\bm{m}_{*}}$ belonging $\mathcal{N}_{2,9(L+1)\log_{2}L}$ with the width $28n^{d+1}(N+d)\log_{2}(8N)$ such that
$\|f(\bm{x})-\gamma_{\bm{m}_{*}}(\bm{x})\|_{W^{2,p}(\Omega_{\bm{m}_{*}})}\leq C_{3}(n,d,a,b)\|f\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2(n-2)/d}L^{-2(n-2)/d}$
$\|f(\bm{x})-\gamma_{\bm{m}_{*}}(\bm{x})\|_{W^{1,p}(\Omega_{\bm{m}_{*}})}\leq C_{3}(n,d,a,b)\|f\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2(n-1)/d}L^{-2(n-1)/d}$
$\|f(\bm{x})-\gamma_{\bm{m}_{*}}(\bm{x})\|_{L^{p}(\Omega_{\bm{m}_{*}})}\leq C_{3}(n,d,a,b)\|f\|_{W^{n,p}(\Omega_{\bm{m}_{*}})}N^{-2n/d}L^{-2n/d},$ (4.23)
where $C_{3}=C_{1}+4^{\frac{2d+p}{2p}}C_{2}$ is the constant independent of $N,L$. The final inequality arises from the fact that $\frac{2n}{d}\leq 2n$ holds for values of $d$ and $n$ greater than or equal to 1.
Similarly, we can construct a $\gamma_{\bm{m}}$ achieving the same order of Eq. (4.23).
Denote a sequence of the neural network to achieve performance in Proposition 4 as $\{s_{\bm{m}}(\bm{x}):=\lambda_{\bm{m},2}\circ\lambda_{\bm{m},1})(\bm{x})\}_{\bm{m}\in\{1,2\}^{d}}$.
Now we define the piece-wise constant function $\phi_{\bm{\alpha}}(\bm{\phi}_{2}(\bm{x}))$ (4.18) defined on $\Omega_{\bm{m}}$ as $p_{\bm{m}}$, then we have
$\gamma(\bm{x})$ $=\sum_{\bm{m}\in\{1,2\}^{d}}s_{\bm{m}}(\bm{x})\cdot\gamma_{\bm{m}}(\bm{x})=\sum_{\bm{m}\in\{1,2\}^{d}}s_{\bm{m}}(\bm{x})\cdot\sum_{|\bm{\alpha}|\leq n-1}p_{\bm{m}}\cdot\bm{x}^{\bm{\alpha}}$
$=\sum_{|\bm{\alpha}|\leq n-1}\sum_{\bm{m}\in\{1,2\}^{d}}s_{\bm{m}}(\bm{x})\cdot p_{\bm{m}}\cdot\bm{x}^{\bm{\alpha}}$
$=\sum_{|\bm{\alpha}|\leq n-1}\sum_{\bm{m}\in\{1,2\}^{d}}(\lambda_{\bm{m},2}\circ\lambda_{\bm{m},1})(\bm{x})\cdot p_{\bm{m}}\cdot\bm{x}^{\bm{\alpha}}$
$=\sum_{|\bm{\alpha}|\leq n-1}\sum_{\bm{m}\in\{1,2\}^{d}}\phi_{6}(\phi_{6}((\lambda_{\bm{m},2},\textbf{Id})\circ(\lambda_{\bm{m},1},p_{\bm{m}})),\bm{x}^{\bm{\alpha}}),$ (4.24)
where we denote $(f_{1},f_{2})\circ(g_{1},g_{2})$ as $(f_{1}\circ g_{1},f_{2}\circ g_{2})$. It is easy to check that $\gamma$ is a DSRN in $\mathcal{N}_{1,9(L+1)\log_{2}4L}$ based on Fig. 5. Furthermore, the width of $\gamma_{\bm{m}_{*}}$ is $2^{d+6}n^{d+1}(N+d)\log_{2}(8N)$.
Figure 5: The structure of $\gamma(\bm{x})$.